OpenStudy (kpop4life123):

I WILL FAN AND MEDAL!!!!!!!!!!!!!!!! There are (35)2 ⋅ 30 leaves on a tree. What is the total number of leaves on the tree?

OpenStudy (anonymous):

35*2=70*30=2100 leaves

OpenStudy (anonymous):

Does that make sense?

OpenStudy (kpop4life123):

its not one of the options

OpenStudy (gabylovesyou):

is (35)2 the same as (35)^2 ?

OpenStudy (kpop4life123):

3^3 3^7 3^10 3^25

OpenStudy (gabylovesyou):

Those are the options ?

OpenStudy (kpop4life123):

oh sorry its (3^5) x 3^0

OpenStudy (gabylovesyou):

3^5 = 3 * 3 * 3 * 3 * 3 = 243 3^0 = 1 243 * 1 = 243

OpenStudy (gabylovesyou):

i dont get ur choices tho.

OpenStudy (kpop4life123):

the question is There are (3^5)^2 ⋅ 3^0 leaves on a tree. What is the total number of leaves on the tree? A)^33 B)3^7 C)3^10 D)3^25

OpenStudy (gabylovesyou):

lol is that the final question ? uve typed it different ways lmao

OpenStudy (kpop4life123):

cuz i corrected it lol

OpenStudy (gabylovesyou):

ok so start with parenthesis. 3^5 = 243 243^2 * 3^0 59049 * 1 from your choices, which one equals 59049 ?

OpenStudy (kpop4life123):

THANK U LOL!!!!!!!!!!!!!!!

OpenStudy (kpop4life123):

I GOT THE ANSWER NOW

OpenStudy (kpop4life123):

THANKS LOL

OpenStudy (gabylovesyou):

3^10 correct ? lol

OpenStudy (gabylovesyou):

your welcome. lol

OpenStudy (kpop4life123):

Yes its correct lol